# RDA-INR: Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations

### v2

@@ Abstract
Diffeomorphic registration frameworks such as Large Deformation Diffeomorphic Metric Mapping (LDDMM) are used in computer graphics and the medical domain for atlas building, statistical latent modeling, and pairwise and groupwise registration. In recent years, researchers have developed neural network-based approaches regarding diffeomorphic registration to improve the accuracy and computational efficiency of traditional methods. In this work, we focus on a limitation of neural network-based atlas building and statistical latent modeling methods, namely that they either are (i) resolution dependent or (ii) disregard any data/problem-specific geometry needed for proper mean-variance analysis. In particular, we overcome this limitation by designing a novel encoder based on resolution-independent implicit neural representations. The encoder achieves resolution invariance for LDDMM-based statistical latent modeling. Additionally, the encoder adds LDDMM Riemannian geometry to resolution-independent deep learning models for statistical latent modeling. We showcase that the Riemannian geometry aspect improves latent modeling and is required for a proper mean-variance analysis. Furthermore, to showcase the benefit of resolution independence for LDDMM-based data variability modeling, we show that our approach outperforms another neural network-based LDDMM latent code model. Our work paves a way to more research into how Riemannian geometry, shape/image analysis, and deep learning can be combined.
Keywords shape space  $\cdot$ Riemannian geometry  $\cdot$ principal geodesic analysis  $\cdot$ LDDMM  $\cdot$ diffeomorphic registration  $\cdot$ latent space  $\cdot$ implicit neural representations
@@ 1 Introduction
Table 1: Table comparing neural network-based diffeomorphic registration methods. The columns showcase the features that a paper treats or does not treat. Our novel RDA-INR framework is the only work that addresses all the features.
@@
Atlas-building Stat. latent modelling Physical consistency Res. indep.
@@
Various registration frameworks are available, among which diffeomorphic registration frameworks such as Stationary Velocity fields (SVF) [22, 23, 24, 25] and Large Deformation Diffeomorphic Metric Mapping (LDDMM). One specific registration task is pairwise registration where a registration between exactly two objects is calculated [26, 27, 28, 29, 30, 31, 1]. To register multiple objects simultaneously, groupwise registration algorithms [32, 33, 34, 30] register all the objects to the same template. In this process, the template can also be estimated to perform atlas building. The estimated atlas is used for proper assessment of the differences between objects. Finally, besides plain diffeomorphic registration, LDDMM is also used for modeling the variability of the data. Specifically, using the LDDMM Riemannian distance, one can use the manifold variant of principal component analysis (PCA), called principal geodesic analysis (PGA), to estimate the factors of variation given the data [35, 36].
Recently, these tasks have been solved via neural network models. For instance, neural network models exist for speeding up the computation of pairwise registrations [2, 3, 4, 5, 11, 12], groupwise registrations [5, 7, 14, 13, 17], and joint groupwise registration and atlas building [7, 14, 17]. Furthermore, attempts have been made to improve LDDMM PGA by adding neural networks [16, 15]. Finally, all the approaches described above require a specific discretization of the objects and the underlying registration domain. Often one does not immediately know the optimal discretization, one might get memory issues when one refines the discretization, and neural network models might not generalize well to other resolutions. Recently, resolution-independent methods avoiding this discretization showed improved performance for diffeomorphic registration [21, 10, 18, 8, 9] and non-diffeomorphic registration [37].
Table 1 compares a comprehensive selection of relevant and representative neural network methods based on their main features. In particular, we focus on atlas building, statistical latent modeling, physical consistency using LDDMM Riemannian geometry, and resolution invariance/independence. In this work, we only classify models as physically consistent if they use LDDMM, as it enables time-dependent non-autonomous behavior. In contrast, SVF, for example, only allows autonomous stationary velocity vector fields. From the table, we notice that many methods do not have all the properties. For instance, to the best of our knowledge, no method exists for performing atlas building and statistical latent modeling that is physically consistent and resolution independent.
There are two types of algorithms for joint atlas building and statistical latent modeling. On the one hand, we have models that consider resolution-independent methods for shape data [38, 39, 21] but do not consider LDDMM physical consistency. This physical consistency is required for properly dealing with longitudinal data and to properly calculate the Fréchet mean and variance of the dataset:
where $d$ is the Riemannian distance induced by LDDMM and $\{O_{i}\}_{i=1}^{N}$ are $N$ images or shapes. On the other hand, we have the LDDMM literature that considers physically consistent methods [16, 15] that are not resolution-independent.
1.1 Contribution
Our work aims to illustrate the general benefit of the LDDMM physical consistency and the benefit of resolution-independence properties. To achieve this goal, we introduce a novel model called Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations (RDA-INR). Our model is designed to deal with the previously mentioned issues of joint atlas building and statistical latent modeling. Although our model is also applicable to images, we mainly focus on the special case of shapes.
Figure 1 shows how we arrive at our model. More precisely, our RDA-INR adds resolution-independent implicit neural representations (INRs) to LDDMM PGA and adds Riemannian geometry to deep learning models for shape variability modeling. Our method shares similarities with the deformable template model developed by Sun et al. [21]. However, our model differs from theirs as we merge it with the LDDMM framework. Specifically, we use the Riemannian deformation cost of LDDMM as a deformation regularizer. Moreover, instead of the signed distance representation of shapes used by Sun et al. [21], we consider point cloud and mesh data.
First, we investigate the benefit of physical consistency. We treat the situation depicted in Figure 2 and show that the Riemannian geometry is required for a proper mean-variance analysis. More precisely, we show that the Riemannian geometry is needed to simultaneously achieve two objectives: obtaining a template as the Fréchet mean of the data and obtaining physically plausible deformations that align the template to the data. Finally, we show that the Riemannian geometry improves the latent modeling as it leads to improved reconstruction generalization and more robustness to noise.
Figure 2: Three latent model classes defined by mean-variance analysis capabilities. The calculation of the Fréchet mean and the variance of a dataset depends on the length of the shortest path between objects. However, not all latent models take this into account, which results in three categories of methods: methods considering the geodesic distance, methods that do not consider the geodesic distance but still obtain a good Fréchet mean, and methods that do not consider the geodesic distance and do not obtain a good Fréchet mean.
While the previous results illustrate the importance of physical consistency, we also treat the importance of resolution independence by comparing our method to DAE [16]. We show that our model obtains higher-quality templates and higher-quality reconstructions, demonstrating the benefit of resolution independence.
@@ 1.2 Outline
We organize the rest of the paper as follows. In Section 2, we discuss the related works. Subsequently, in Section 3, we treat the background theory needed for our method. Specifically, we treat INRs and the implicit representation of shapes (Section 3.1), the LDDMM framework (Section 3.2), and LDDMM PGA (Section 3.3). Section 4 introduces our neural network model. In Section 5 we compare this model to the same model with a non-Riemannian regularization. We also compare our method to DAE [16]. Finally, we summarize the paper, discuss some limitations, and discuss future work in Section 6. In the appendices, we provide additional experiments and details, and some basics of Riemannian geometry.
@@ 2.1 Principal geodesic analysis
Bone et al. [16] and Hinkle et al. [15] construct an autoencoder that resembles LDDMM PGA [35]. While Hinkle et al. [15] is only applicable to images, Bone et al. [16] is also applicable to meshes. Similar to our work, both use an ordinary differential equation (ODE) to deform a learned template into a reconstruction. Moreover, they introduce a regularizer on the velocity vector fields to let the ODE flow be a diffeomorphism. In contrast, while they use a specific resolution for the images, shapes, and diffeomorphisms, we construct a resolution-independent method. For instance, while Bone et al. [16] consider the mesh representation of shapes, we consider the state-of-the-art implicit shape representation. Implicit shape representations allow us to acquire infinite-resolution shape reconstructions because the shapes are represented by the zero level set of a function.
@@ 2.2 Riemannian geometry for latent space models
Latent space models based on neural networks are used for, among others, dimensionality reduction, clustering, and data augmentation. Most of these models have the drawback that distances between latent codes do not always correspond to a ’semantic’ distance between the objects they represent. One mathematical tool to solve this issue is Riemannian geometry. Arvanitidis et al. [51] design a Riemannian distance between latent codes by defining the corresponding Riemannian metric. This metric is defined as the standard Euclidean metric pulled back under the action of the decoder. The shortest paths under this Riemannian metric are found by solving the geodesic equation via a numerical solver. However, one can also find the shortest path by solving the geodesic equation via Gaussian processes and fixed point iterations [52]. Another approach is to solve the geodesic minimization problem [53, 54]. In Arvanitidis et al. [55] a surrogate Riemannian metric is proposed that approximates the Euclidean pullback metric and yields more robust calculations of shortest paths. In Chadebec et al. [54], the Riemannian metric is incorporated into a Hamiltonian Markov Chain to sample from the posterior distribution of a VAE. Geng et al. [56] enforce a Riemannian distance structure by encouraging Euclidean distances in latent space to equal the Riemannian distance in the original space. Consequently, linear interpolations correspond to points on the learned manifold.
Besides using the Riemannian metrics and the shortest paths for defining good distances, creating good interpolations, and defining probability distributions on the latent space, they are also used for improved shape generation [54], for improved clustering [57, 54], and for data augmentation [58].
Our novel latent model that uses implicit neural representations (INRs) is related to the Riemannian geometry for latent space models literature. The reason is that we embed a Riemannian distance between a learned template and the other reconstructed data points by using LDDMM as the Riemannian distance. The most relevant work that also attempts to incorporate shape Riemannian geometry into an INR latent space model is Atzmon et al. [59]. However, their model can not perform shape encoding and shape registration jointly. Moreover, it does not enforce a diffeomorphic structure.
2.3 Neural ODEs
The neural ODE (NODE) [60] is a first-order differential equation where the time derivative is parameterized by a neural network. It has applications in learning dynamics [61, 62, 63, 64], control [65, 66], generative modeling [60, 67], and joint shape encoding, reconstruction, and registration [68, 20, 21]. Furthermore, there is a bidirectional connection between the NODE and optimal transport (OT). First, due to the fluid dynamics formulation of OT [69], NODEs can be used to solve high-dimensional OT problems [70]. In addition, the learned dynamics can be complex and require many time steps to be solved accurately. As a consequence, training NODEs can be challenging and time-consuming. OT-inspired regularization functionals solve this issue as they simplify the dynamics and reduce the time needed to solve the NODE [71, 72].
@@ 3.1 Implicit neural representations and implicit shape representations
In many variational problems, we need to optimize for an unknown image, a solution to a partial differential equation, a velocity vector field for registration, or a shape. The idea behind INRs is to represent such objects by a neural network $f_{\theta}:\Omega\rightarrow\mathbb{R}^{n}$, where $\Omega$ can be a spatial or a spatiotemporal domain. This representation avoids discretization of the objects and by optimizing over the neural network weights, we obtain a solution in an infinite dimensional space. For example, methods that use $f_{\theta}$ as parameterization of a velocity vector field or a deformation field achieve good performance in diffeomorphic [10, 18, 8, 9] and non-diffeomorphic registration [37].
When representing shapes with INRs, the implicit shape representation is of interest. A shape $\mathcal{S}:=\{x\mid f(x)=c\}$ is described by the $c$ level set of a function $f:\Omega\rightarrow\mathbb{R}$ on the image domain $\Omega\subset\mathbb{R}^{d}$. Given a shape $\mathcal{S}=\partial\mathcal{S}_{\text{int}}$ with interior $\mathcal{S}_{\text{int}}\subset\Omega$, two examples of common implicit representations are the signed distance function (SDF) with $c=0$ and the occupancy function (OCC) with $c=0.5$, respectively:
$\text{SDF}_{\mathcal{S}}(x)$ =$\begin{cases}\inf_{y\in\mathcal{S}}\lVert x-y\rVert_{2}\quad\quad\text{if }x\in\mathcal{S}_{\text{int}},\\ -\inf_{y\in\mathcal{S}}\lVert x-y\rVert_{2}\quad\text{if }x\notin\mathcal{S}_{\text{int}},\\ \end{cases}\end{aligned}$ (1)
$\text{OCC}_{\mathcal{S}}(x)$ =$\begin{cases}0\qquad\qquad\qquad\qquad\text{if }x\notin\mathcal{S}_{\text{int}},\\ 0.5\qquad\qquad\qquad\quad\,\text{if }x\in\mathcal{S},\\ 1\qquad\qquad\qquad\qquad\text{if }x\in\mathcal{S}_{\text{int}}.\\ \end{cases}\end{aligned}$ (2)
Unlike the traditional discrete representations such as point clouds, meshes, or voxel grids, the implicit shape representation is continuous as it represents the shape using a function. Consequently, we can obtain a mesh or point cloud at an arbitrary resolution by using marching cubes [73]. INR methods that use this continuous representation, such as DeepSDF [74], SIREN [75], and Occupancy Network [76], yield state-of-the-art shape encodings and reconstructions. However, unlike the discrete representations, the implicit representation has the disadvantage of not providing point correspondences between shape deformations. To remedy this issue, recent works [38, 39, 21] create template-based INR models that allow for joint registration and encoding. These models parameterize a template shape with an INR and transform it into other shapes to allow for shape matching.

### v3

@@ Abstract
Diffeomorphic registration frameworks such as Large Deformation Diffeomorphic Metric Mapping (LDDMM) are used in computer graphics and the medical domain for atlas building, statistical latent modeling, and pairwise and groupwise registration. In recent years, researchers have developed neural network-based approaches regarding diffeomorphic registration to improve the accuracy and computational efficiency of traditional methods. In this work, we focus on a limitation of neural network-based atlas building and statistical latent modeling methods, namely that they either are (i) resolution dependent or (ii) disregard any data- or problem-specific geometry needed for proper mean-variance analysis. In particular, we overcome this limitation by designing a novel encoder based on resolution-independent implicit neural representations. The encoder achieves resolution invariance for LDDMM-based statistical latent modeling. Additionally, the encoder adds LDDMM Riemannian geometry to resolution-independent deep learning models for statistical latent modeling. We investigate how the Riemannian geometry improves latent modeling and is required for a proper mean-variance analysis. To highlight the benefit of resolution independence for LDDMM-based data variability modeling, we show that our approach outperforms current neural network-based LDDMM latent code models. Our work paves the way for more research into how Riemannian geometry, shape respectively image analysis, and deep learning can be combined.
keywords:
shape space, Riemannian geometry, principal geodesic analysis, LDDMM, diffeomorphic registration, latent space, implicit neural representations
{MSCcodes}
58D05, 53A05, 58B20, 68T07, 62-07, 68U10
@@ 1 Introduction
Various diffeomorphic registration frameworks are available in the literature, for instance, Stationary Velocity fields (SVF) [3, 7, 69] or Large Deformation Diffeomorphic Metric Mapping (LDDMM). One specific registration task is pairwise registration where a registration between exactly two objects is calculated [11, 67, 50, 70, 80, 82, 2]. To register multiple objects simultaneously, groupwise registration algorithms [40, 46, 47, 80] register all the objects to the same template. In this process, the template can also be estimated to perform atlas building. The estimated atlas is used for proper assessment of the differences between objects. Finally, besides plain diffeomorphic registration, LDDMM is also used for modeling the variability of the data. Specifically, using the LDDMM Riemannian distance, one can use the manifold variant of principal component analysis (PCA), called principal geodesic analysis (PGA), to estimate the factors of variation given the data [81, 84]. The Lagrangian nature of LDDMM acts as a physical modeling basis (physical consistency) providing access to Fréchet mean and variance. More precisely, we use the term physical consistency as LDDMM’s geodesic structure enables time-dependent non-autonomous behavior and is reminiscent of Lagrangian structures used in physics via the least-action principle.
Recently, the different registration tasks have also been addressed via neural network methods. Table 1, positioned near the end of the introduction, presents and compares a comprehensive selection of them. Specifically, there are neural network models that speed up the computation of pairwise registrations [39, 43, 60, 78, 38, 71], groupwise registrations [78, 24, 26, 52, 72], and joint groupwise registration and atlas building [24, 26, 72]. Furthermore, attempts have been made to improve LDDMM PGA by adding neural networks [14, 36]. Finally, all the approaches described above require a specific spatial discretization of the objects and the underlying registration domain. Often one does not a-priori know the optimal discretization, one might get memory issues when one refines the discretization, and neural network models might not generalize well to other resolutions. Recently, resolution-independent methods avoiding this spatial discretization showed improved performance for diffeomorphic registration [63, 86, 75, 33, 65] and non-diffeomorphic registration [73]. This motivates the need for having resolution-independent methods for LDDMM PGA, which achieves joint atlas building and statistical latent modeling.
In the literature, there are two types of approaches for joint atlas building and statistical latent modeling. On the one hand, we have the aforementioned LDDMM literature that considers methods that have LDDMM physical consistency [14, 36] but are not resolution-independent. On the other hand, we have models that consider resolution-independent implicit neural representation (INR) methods for shape data [25, 85, 63] but do not consider LDDMM physical consistency. This physical consistency is required for properly dealing with longitudinal data and to properly calculate the Fréchet mean and variance of the dataset:
where $d$ is the Riemannian distance induced by LDDMM and $\{O_{i}\}_{i=1}^{N}$ are $N$ images or shapes. When the model does not consider this Riemannian structure, one might not calculate the actual geodesics, which are needed for calculating a proper estimate of the variance. Moreover, the estimated mean might not be consistent with the data, which is detrimental if you want to study a dataset via a sensible average or atlas.
1.1 Contributions
Our work addresses the previously mentioned issues of joint atlas building and statistical latent modeling to study the benefits of (i) LDDMM physical consistency and (ii) resolution-independence properties. Specifically, we connect the shape INR literature with the LDDMM PGA literature, where the latter focuses on image data. To have a general viewpoint, we address this connection mainly from a point cloud or mesh perspective instead of a segmentation mask perspective. Our main contributions are:
  1. 1.
    We introduce a novel model called Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations (RDA-INR) that deals with the aforementioned issues. Figure 1 shows how our model is designed. Specifically, our RDA-INR adds:
    • Riemannian geometry to the INR literature on shape variability modeling by using LDDMM PGA and its Riemannian regularization functional,
    • resolution-independent INRs to LDDMM PGA.
  2. 2.
    Our latent model is the first in the registration literature using image-based (implicit neural) data representations for dealing directly with point cloud or mesh data.
  3. 3.
    We show that the Riemannian regularization functional is required for a proper mean-variance analysis, i.e. avoiding the middle and right category in Figure 2. More precisely, the Riemannian geometry is needed to simultaneously obtain: a template as Fréchet mean of the data and physically plausible deformations aligning the template to the data.
  4. 4.
    We show that the Riemannian regularization functional can improve the latent modeling compared to a non-Riemannian regularization functional. More precisely, it can lead to improved reconstruction generalization and more robustness to noise.
  5. 5.
    By comparing our model to a mesh-based neural network latent model from the LDDMM literature, we show that resolution independence yields higher-quality templates and higher-quality reconstructions.
Figure 2: Latent model classes for mean-variance analysis. The calculation of the Fréchet mean and the variance of a dataset depend on the length of the geodesics between objects. However, not all latent models take this into account, which results in three categories of methods: methods considering geodesics and obtaining a data-consistent Fréchet mean, methods that do not consider geodesics but still obtain a Fréchet mean consistent with the data, and methods that do not obtain such a Fréchet mean.
@@ 1.1 Contributions
Table 1: Neural network-based diffeomorphic registration methods. The columns show the features of the papers: atlas building, statistical latent modeling, physical consistency using LDDMM Riemannian geometry, and resolution invariance. Our novel RDA-INR framework is the only work that addresses all the features.
@@
Atlas building Stat. lat. modeling Riemannian Res. indep.
@@ 1.2 Outline
We organize the rest of the paper as follows. In Section 2, we discuss the related works. Subsequently, in Section 3, we treat the background theory needed for our method. Specifically, we treat INRs and the implicit representation of shapes (Section 3.1), the LDDMM framework (Section 3.2), and LDDMM PGA (Section 3.3). Section 4 introduces our neural network model. In Section 5 we compare this model to the same model with a non-Riemannian regularization. We also compare our method to DAE [14]. Finally, we summarize the paper, discuss some limitations, and discuss future work in Section 6. In the appendices, we provide additional experiments and details, and some basics of Riemannian geometry. Specifically, the appendices present an interesting result: our Riemannian regularization functional is novel in the LDDMM literature, yet still related to the standard LDDMM regularizers.
@@ 2.1 Principal geodesic analysis
Bone et al. [14] and Hinkle et al. [36] construct an autoencoder that resembles LDDMM PGA [81]. While Hinkle et al. [36] is only applicable to images, Bone et al. [14] is also applicable to meshes. Similar to our work, both use an ordinary differential equation (ODE) to deform a learned template into a reconstruction. Moreover, they introduce a regularizer on the velocity vector fields to let the ODE flow be a diffeomorphism. In contrast, while they use a specific resolution for the images, shapes, and diffeomorphisms, we construct a resolution-independent method. For instance, while Bone et al. [14] consider the mesh representation of shapes, we consider the state-of-the-art implicit shape representation. Combining this representation with INRs allows us to acquire a resolution-independent approach where we obtain a mesh or point cloud at an arbitrary resolution by applying marching cubes [44].
@@ 2.2 Riemannian geometry for latent space models
Latent space models based on neural networks are used for, among others, dimensionality reduction, clustering, and data augmentation. Most of these models have the drawback that distances between latent codes do not always correspond to a ’semantic’ distance between the objects they represent. One mathematical tool to solve this issue is Riemannian geometry. Arvanitidis et al. [5, 4] design a Riemannian distance between latent codes by defining a Riemannian metric. This metric is the standard Euclidean metric pulled back under the action of the decoder or a surrogate metric approximating the pullback. Shortest paths under these Riemannian metrics are found by solving the geodesic equation [5, 6] or the geodesic minimization problem [19, 16]. The resulting paths yield good distances and create good interpolations. Furthermore, the Riemannian metric can be used for improved clustering [77, 16] and for improved generative modeling. In particular, Chadebec et al. [16] incorporate the Riemannian metric into a Hamiltonian Markov Chain that samples from a variational autoencoder’s posterior distribution.
Our novel latent model that uses implicit neural representations (INRs) is related to the Riemannian geometry for latent space models literature. Although the Euclidean distances in our latent space do not necessarily approximate LDDMM distances, we do embed a Riemannian distance between a learned template and the other reconstructed data points by using LDDMM as the Riemannian distance. The most relevant work that also attempts to incorporate shape Riemannian geometry into an INR latent space model is Atzmon et al. [8]. However, their model can not perform shape encoding and shape registration jointly. Moreover, it does not enforce a diffeomorphic structure.
2.3 Neural dynamics
The neural ODE (NODE) [20] is a first-order differential equation where the time derivative is parameterized by a neural network. It has applications in learning dynamics [30, 23, 66], control [56, 74], generative modeling [20, 76], and joint shape encoding, reconstruction, and registration [53, 32, 63]. Furthermore, there is a bidirectional connection between the NODE and optimal transport (OT). First, due to the fluid dynamics formulation of OT [12], NODEs can be used to solve high-dimensional OT problems [58]. In addition, the learned dynamics can be complex and require many time steps to be solved accurately. As a consequence, training NODEs can be challenging and time-consuming. OT-inspired regularization functionals solve this issue as they simplify the dynamics and reduce the time needed to solve the NODE [28, 55].
@@ 3.1 Implicit neural representations and implicit shape representations
In many variational problems, we optimize for an unknown image, a solution to a partial differential equation, a velocity vector field for registration, or a shape. The idea behind an implicit neural representation (INR) is to represent such objects by a neural network $f_{\theta}\colon\Omega\rightarrow\mathbb{R}^{n}$, where $\Omega$ can be a spatial or a spatiotemporal domain. In contrast to, for instance, finite element solutions, which also lie in an infinite dimensional space, the neural network representation avoids a spatial or spatiotemporal discretization of the objects. Furthermore, INRs yield a solution in an infinite dimensional space by optimizing over elements from the finite-dimensional submanifold defined by the neural network weights. Methods employing this approach and using $f_{\theta}$ as parameterization of a velocity vector field or a deformation field, achieve good performance in diffeomorphic [86, 75, 33, 65] and non-diffeomorphic registration [73].
When representing shapes with INRs, the implicit shape representation is of interest. A shape $\mathcal{S}:=\{x\mid f(x)=c\}$ is described by an image via the $c$ level set of an image function $f\colon\Omega\rightarrow\mathbb{R}$ on the domain $\Omega\subset\mathbb{R}^{d}$. Given a shape $\mathcal{S}=\partial\mathcal{S}_{\text{int}}$ with a bounded open set $\mathcal{S}_{\text{int}}\subset\Omega$ as interior, two examples of common implicit representations are the signed distance function (SDF) with $c=0$ and the occupancy function (OCC) with $c=0.5$, respectively:
(1) $\text{SDF}_{\mathcal{S}}(x)$ =$\begin{cases}\inf_{y\in\mathcal{S}}\left\lVert x-y\right\rVert_{2}\quad\quad\text{if }x\in\mathcal{S}_{\text{int}},\\ -\inf_{y\in\mathcal{S}}\left\lVert x-y\right\rVert_{2}\quad\text{if }x\notin\mathcal{S}_{\text{int}},\\ \end{cases}\end{aligned}$
(2) $\text{OCC}_{\mathcal{S}}(x)$ =$\begin{cases}0\qquad\qquad\qquad\qquad\text{if }x\in\Omega\backslash\overline{\mathcal{S}_{\text{int}}},\\ 0.5\qquad\qquad\qquad\quad\,\text{if }x\in\mathcal{S},\\ 1\qquad\qquad\qquad\qquad\text{if }x\in\mathcal{S}_{\text{int}}.\\ \end{cases}\end{aligned}$
Unlike the traditional methods that use point clouds, meshes, or voxel grids, INR methods that use the implicit representation, such as DeepSDF [57], SIREN [61], and Occupancy Network [48], are spatial discretization-independent. Furthermore, they yield state-of-the-art shape encodings and reconstructions, which are obtained by extracting a level set at arbitrary resolution, for instance, via marching cubes [44]. However, unlike deforming meshes and point clouds by deforming their vertices, morphing an implicit representation does not immediately create point correspondences between the initial and deformed shapes. To remedy this issue, recent works [25, 85, 63] create template-based INR models that allow for joint registration and encoding. These models parameterize a template shape with an INR and transform it into implicit representations of other shapes via image registration techniques. Subsequently, the model uses the image registrations to the template INR to perform shape matching.